An organic compound X contains 40% carbon, 6.7% hydrogen and the rest oxygen. The vapour density of X is 30. The compound reacted with a compound Y and a colourless gas Z which turn limewater milky. On boiling under reflux, the mixture of X and W, gives a sweet smelling liquid G was formed.i. determine the empirical and molecular formula of X.ii. identify Y, Z, W and G.iii. Give the formula of the functional group present in Xiv. Deduce the structure of the compound Xv. Name the reaction which take place between X and W, and write the equation for the reaction of X and W.

1
Expert's answer
2022-06-09T16:43:03-0400
  • The molar mass of X

To find the molar mass of compound X we can use the formula M=DxM(H2) :

[Molar mass of compound X M(X)] = [vapour density D(X)]x[molar mass of hydrogen M(H2)]

M(X) = 30 x 2 g/mol = 60 g/mol

  • The formula of compound X

a) 40% of carbon in compound X is 0.4 x 60 g/mol = 24 g/mol

To determine number of carbom atoms in each X molecule it is needed to divide 24g/mol to molar mass of carbon or 12 g/mol:

N(C) = 24 g/mol : 12 g/mol = 2 

b) 6.7% of hydrogen in compound X is 0.067 x 60 g/mol = 4 g/mol

N(H) = 4 g/mol : 1 g/mol = 4

c) Compound X contains only C, H, and O. Percentage of oxygen is 100% - 40% - 6.7% = 53.3%.

53.3% of oxygen in compound X is 0.533 x 60 g/mol = 32 g/mol .

N(O) = 32 g/mol : 16 g/mol = 2.

The molecular formula of X is C2H4O2, the empirical formula is CH2O.

  • Limewater 

Limewater is a saturated solution of calcium hydroxide. It is turned milky because of the reaction Ca(OH)2 with CO2: Ca(OH)2(aq) + CO2(g) = CaCO3(s) + H2O(l)

It means that a colourless gas Z is carbon dioxide or CO2. It can be obtained by reaction of NaHCO3 (compound Y) with an acid. The only possible acid with a molecular formula of C2H4O2 is acetic acid (compound X). The reaction between compounds X and Y: CH3COOH(l) + NaHCO3(s) = CH3COONa + CO2(g) + H2O(l)

  • A sweet-smelling liquid

During boiling under reflux with alcohols, acetic acid can form esters. One of the most famous liquid sweet-smelling esters is ethyl acetate CH3COOCH2CH3. It can be obtained by the reaction of acetic acid with ethyl alcohol (compound W):

CH3COOH(l) + CH3CH2OH(l) "\to" CH3COOCH2CH3(l) + H2O(l)


Answers:

i. The empirical formula of X is CH2O, and the molecular formula of X is C2H4O2

ii. Y is NaHCO3, Z is CO2, W is CH3CH2OH

iii. functional group of X is carboxylic acid: -COOH